OpenStudy (anonymous):

I give medals for correct answers! what is the slope-intercept for the points (-1,4) (-5,2)

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

ok thats the slope, but what about the rest?

OpenStudy (anonymous):

IDK! :'(

OpenStudy (anonymous):

\[(y-4)/(x- -1) = (4-2)/ (-1 - -5) =2/4= 1/2\] \[2(y-4)= x+1 \rightarrow y= x/2 + 9\ \rightarrow y=mx+c\] Hence, slope= 1/2, intercept= 9

OpenStudy (anonymous):

thanks!